# revealing an in-context neural scaling law

### v2

@@ Abstract
Pretrained large language models (LLMs) are surprisingly effective at performing zero-shot tasks, including time-series forecasting. However, understanding the mechanisms behind such capabilities remains highly challenging due to the complexity of the models. We study LLMs’ ability to extrapolate the behavior of dynamical systems whose evolution is governed by principles of physical interest. Our results show that LLaMA 2, a language model trained primarily on texts, achieves accurate predictions of dynamical system time series without fine-tuning or prompt engineering. Moreover, the accuracy of the learned physical rules increases with the length of the input context window, revealing an in-context version of neural scaling law. Along the way, we present a flexible and efficient algorithm for extracting probability density functions of multi-digit numbers directly from LLMs.
@@ 1 Introduction
Since the introduction of the transformer architecture (Vaswani et al., 2017), Large language models (LLMs) have shown a variety of unexpected emergent properties, such as program execution (Nye et al., 2021) and multi-step reasoning (Cobbe et al., 2021; Wei et al., 2022; Suzgun et al., 2022).
Despite the empirical success of LLMs, how they compress vast amounts of information and implement complex algorithms within their architecture is not readily discoverable. To this end, recent studies used representation probes to decipher how concepts and functions are encoded in the layers of trained neural networks (Akyürek et al., 2022; Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Park et al., 2023; Hendel et al., 2023).
@@
Our work explores LLMs’ ability to model the world by proposing a new perspective and empirical approach. Inspired by the recent observations that LLMs are capable of in-context time series extrapolation without specific prompting or fine-tuning (Gruver et al., 2023; Jin et al., 2023a), we aim to quantify LLMs’ ability to extrapolate stochastic dynamical systems. We find that, as the number of observed time steps increases, an LLM’s statistical prediction consistently converges to the ground truth transition rules underlying the system; leading to an empirical scaling law, as observed in Figure 1.
@@ 2 Background and related work
In-context learning refers to LLM’s emergent ability to learn from examples included in the prompt (Brown et al., 2020). One example of in-context learning is zero-shot time series forecasting (Gruver et al., 2023).
The work of (Gruver et al., 2023) aims to forecast empirical time series and introduces a tokenization procedure to convert a sequence of floating point numbers into appropriate textual prompts for LLMs. This led to several subsequent studies on the application of LLMs for time series forecasting (Chen et al., 2023; Jin et al., 2023b, a; Dooley et al., 2023; Schoenegger & Park, 2023; Wang et al., 2023; Xu et al., 2023).
@@ 3.1 Prompt generation
Markov processes. Most of our testing data may be modeled as discrete-time Markov chains, where the probability distribution function (PDF) of the next state at time $t+1$ depends solely on the previous state $x_{t}$ at time $t$:
@@
Time series tokenization. An input time series typically consists of ($\sim 10^{3}$) time steps, each represented as a real number. We first rescale each number and represent it using $n$ digits (typically, $n=3$). Each time series is rescaled to the interval $[1.50,8.50]$ so that the number of digits never changes throughout the series. We then follow the scheme introduced in (Gruver et al., 2023) to serialize the time series as strings and tokenize them.
@@ 3.2 Extraction of transition rules
Discrete state space. When the Markov process is discrete and has a finite state space, each state can be represented by a single token. We employ tokens corresponding to the ten number strings: $0,\ldots,9$. We find that even the most sophisticated LLaMA model (LLaMA-70b) can only learn up to 9 discrete states. Therefore, we do not attempt to go beyond 9 distinct states by extending to non-number tokens (see Section A.3.3 for more details).
@@
Figure 2 illustrates our framework for learning discrete Markov chains with LLMs. First, we randomly sample an $n\times n$ transition matrix $(P_{ij})$. We then generate a Markov chain according to $P_{ij}$, tokenize the time series and pass to an LLM with no additional “prompt engineering”. The length of the series is chosen such that the tokenized representation does not exceed the length of the LLM’s context window. We extract the LLM’s prediction for the next state by performing a softmax operation on the output logits corresponding to the $n$ allowed states and discarding all other logits.
Continuous state space. Many stochastic processes, such as the Brownian motion (Einstein, 1905; Perrin, 1909), are supported on continuous state space. For these processes, we represent the value of each state as a multi-digit number and separate each state using the comma symbol “,”. As observed in (Jin et al., 2023a), an LLM prediction of multi-digit values can be naturally interpreted as a hierarchical softmax distribution (Mnih & Hinton, 2008; Challu et al., 2023).
@@
Hierarchy-PDF algorithm. While a single pass through the LLM yields a discretized PDF represented by bins of various widths, we can refine the PDF by querying each coarse bin. For example, to furnish a maximal resolution PDF of a 3-digit number, we need to query all $10^{2}$ combinations of the first two digits of that number. Suppose a time series consists of $S$ values (steps), each represented as $n$ digits. Obtaining a maximal resolution PDF for each value of the entire sequence requires $10^{n-1}S$ forward passes of the LLM. This daunting process could be significantly simplified because most of the $10^{n-1}S$ inputs differ only in the last tokens, and thus one can recursively cache the key and value matrices associated with the shared tokens. The computation can be further reduced by refining only the high-probability bins near the mode, which dominate the loss functions, as shown in Figure 3. Algorithm 1 outlines the Hierarchy-PDF algorithm used to recursively refine the PDF associated with a multi-digit value in a time series (more details available in  Appendix A).
@@ 4 Experiments and Results
This section reports empirical in-context learning results on two example systems: discrete Markov chain and stochastic logistic map. We defer discussion of other systems, reported in Figure 1, to Sections A.3 and A.4. The experiments are repeated ten times with trajectories initiated by different random seeds.
Model choice. All numerical experiments reported in this section are performed using the open-source LLaMA-13b model. While we observe that larger language models, such as LLaMA-70b, may achieve lower in-context loss on some dynamical systems (Section A.3.3), they do not display qualitative differences and affect our conclusions.
@@ 4.1 Markov chains with discrete states
Using the testing procedures elaborated in Section 3.2, we generate 10 Markov chains, each from a distinct and randomly generated transition matrix of size $n=4$.
@@ 4.2 Noisy logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behaviors (Strogatz, 2015). It is governed by the following iterative equation:
@@
where $r\in[1,4)$ is a parameter. The logistic map system becomes stochastic when one introduces small Gaussian perturbations of variance at each step, resulting in modified iterative equation:
@@
The first derivative of $f$ measures how sensitive the local dynamics are to external perturbations. This intuitively explains why the standard deviation of the conditional distribution should be proportional to $f^{\prime}$. We note that the approximation in Equation 1 assumes a small perturbation compared to the second derivative, that is $\sigma^{2}\ll 1/f^{\prime\prime}(x)$.
@@
We again observe a power-law-like decay of the in-context loss function with respect to the length of the observed time series in Figure 5. To achieve low in-context loss, the LLM must learn to predict not only the mean, but also the variance of future steps. This is shown in Figure 6 and discussed further in Section A.9.
@@ 5 Discussion and conclusion
Data leakage. The possibility that LLMs’ accurate predictions of the next time step value are merely due to memorization seems extremely unlikely. A sequence of even one thousand numerical values, encoded with three digits, may cover $10^{3000}$ particular instances, which is well beyond the $\sim 10^{12}$ tokens of the training corpus (Touvron et al., 2023).
@@
In-context neural scaling law. Neural scaling laws (Kaplan et al., 2020) are power laws that characterize how the loss of trained neural networks vary with respect to parameters of the model, such as model size, dataset size, and computational resources. To the best of our knowledge, neural scaling laws have so far only been observed in the training procedure, which updates the weights of neural networks using an explicit algorithm, such as stochastic gradient descent and Adam (Kingma & Ba, 2014). The loss curves observed in the different numerical experiments (see Figures 1 and 18) reveal an additional in-context scaling law for LLMs zero-shot learning of dynamical systems. Further analysis of these scaling laws are resented in Section A.6.
Main conclusions. We showed that, with increasing cardinality in tokenization, and given sufficient context, LLMs can accurately extrapolate not just deterministic trajectories (Section A.4), but also chaotic and stochastic sequences governed by specific transition rules. In the latter case, the extrapolation accuracy is measured in a statistical sense. This suggests that large language models can in-context learn dynamical systems’ time series and predict future states in a manner that maintains fidelity with the underlying principles governing the system’s evolution. Moreover, this behavior is observed without the use of any fine-tuning or “prompt engineering”. Our results suggest that a transformer, trained primarily on textual data, can extract governing principles of numerical sequences observed in-context.
Future directions. The in-context neural scaling law hints at a learning algorithm that LLMs implicitly implement during inference, such as gradient descent (Von Oswald et al., 2023). Characterizing such an algorithm is an open question of broad interest  (Shen et al., 2023). Another exciting future direction is to study the generalization of the observed in-context neural scaling laws for other LLMs, such as GPT4 (OpenAI, 2023), and the newly proposed state space models (Gu & Dao, 2023).
Software and Data
The code and data supporting this study will be deposited on a public GitHub repository upon acceptance.
@@ A.1 Loss Functions
Bhattacharyya distance. For stochastic time series, we use the Bhattacharyya distance to characterize the distance between learned and ground truth transition functions. The Bhattacharyya distance between $P$ and $\tilde{P}$, on a domain $\mathcal{X}$ is defined as (Bhattacharyya, 1943, 1946; Kailath, 1967):
@@
and has been widely employed by feature selection and signal extraction methods (Choi & Lee, 2003; Kailath, 1967). Since $\tilde{P}(X_{t+1}|X_{t})$ takes the form of a hierarchical PDF, one may approximate the integral in Equation 2 via a discrete quadrature rule as
@@
Squared deviations from the mean. For deterministic systems, the true transition functions become delta-functions. As a result, the discretized Bhattacharyya distance from Equation 3 reduces to (see Equation 5 in Section A.2)
@@
which is proportional to the negative log-likelihood (NLL) assigned to the true data by the LLM, plus a constant $C$ [^2] This constant is determined by the base B of the system, and the number of digits n as $C=-\ln\Delta x=n\log B$.. NLL references only the finest bins in the hierarchical PDF and is thus unstable as in-context loss. As an alternative, we use the squared deviations from the mean (SDM) (Kobayashi & Salam, 2000) as the in-context loss for deterministic systems:
@@ A.2 Additional loss functions
KL-divergence. The KL-divergence between two PDFs, $P$ and $\tilde{P}$, is defined as
@@
Although commonly used as the training loss for a variety of machine learning systems, this loss function may suffer from numerical instabilities as the learned transition function $\tilde{P}$ are often close to zero, as shown in Figures 13 and 8, where the probability density is concentrated in small regions of the support.
Discretized Bhattacharyya distance for deterministic systems. For deterministic systems, the ground truth transition function is a delta function. Therefore, the Bhattacharyya distance between it and the Hierarchy-PDF prediction only references the finest bin associated with the true value $x_{\mathrm{true}}$.
@@
Higher moments and kurtosis. While the Bhattacharyya distance and SDM measure the agreement between the extracted transition rules $\tilde{P}$ and the ground truth distribution $P$, they do not explicitly characterize the type of the distribution (e.g., Gaussian or uniform). We employ the kurtosis as an additional measure to assess whether the LLM recovers the correct shape of $P$. The kurtosis of a distribution $P$ is defined as (Joanes & Gill, 1998)
@@ A.3.1 Brownian motion
Brownian motion is an example of a continuous-time stochastic process (Einstein, 1905), and is described by a stochastic differential equation (SDE):
@@
where $X_{t}$ represents the state of the system at time $t$, $\mu$ is the drift coefficient, $\sigma$ is the volatility coefficient, and $dW_{t}$ is the increments of a Wiener process (Revuz & Yor, 2013), modeling the randomness of motion.
@@
To simulate trajectories of Brownian motion, we use the Euler–Maruyama method (Platen, 1999), which discretizes Equation 8 as $X_{t+\Delta t}=X_{t}+\mu\Delta t+\sigma\sqrt{\Delta t}Z$, where $\Delta t$ is the time resolution, and $Z\sim\mathcal{N}(0,1)$ is a random variable that follows a standard Gaussian distribution. The Euler–Maruyama method may also be written as a conditional distribution:

### v3

# LLMs learn governing principles of dynamical systems, revealing an in-context neural scaling law

@@ Abstract
We study LLMs’ ability to extrapolate the behavior of various dynamical systems, including stochastic, chaotic, continuous, and discrete systems, whose evolution is governed by principles of physical interest. Our results show that LLaMA-2, a language model trained on text, achieves accurate predictions of dynamical system time series without fine-tuning or prompt engineering. Moreover, the accuracy of the learned physical rules increases with the length of the input context window, revealing an in-context version of a neural scaling law. Along the way, we present a flexible and efficient algorithm for extracting probability density functions of multi-digit numbers directly from LLMs. [^1] Our code is publicly available at https://github.com/AntonioLiu97/llmICL.
LLMs learn governing principles of dynamical systems, revealing an in-context neural scaling law
[…]
@@ 1 Introduction
Since the introduction of the transformer architecture Vaswani et al. (2017), Large language models (LLMs) have shown a variety of unexpected emergent properties, such as program execution Nye et al. (2021) and multi-step reasoning Suzgun et al. (2022).
Our work explores LLMs’ ability to model the world by proposing a new perspective and empirical approach. Inspired by the recent observations that LLMs are capable of in-context time series extrapolation without specific prompting or fine-tuning Gruver et al. (2023); Jin et al. (2023a), we aim to quantify LLMs’ ability to extrapolate stochastic dynamical systems. We find that, as the number of observed time steps increases, an LLM’s statistical prediction consistently converges to the ground truth transition rules underlying the system; leading to an empirical scaling law, as observed in Figure 1.
@@ 2 Background and related work
In-context learning refers to LLM’s emergent ability to learn from examples included in the prompt Brown et al. (2020). One example of in-context learning is zero-shot time series forecasting Gruver et al. (2023). This work aims to forecast empirical time series and introduces a tokenization procedure to convert a sequence of floating point numbers into appropriate textual prompts for LLMs. This led to several subsequent studies on the application of LLMs for time series forecasting Chen et al. (2023); Jin et al. (2023b, a); Dooley et al. (2023); Schoenegger and Park (2023); Wang et al. (2023); Xu et al. (2023).
@@ 3.1 Prompt generation
Markov processes.
Most of our testing data may be modeled as discrete-time Markov chains, where the probability density function (PDF) of the next state at time $t+1$ depends solely on the previous state, $x_{t}$, at time $t$:
@@
Time series tokenization.
An input time series typically consists of ($\sim 10^{3}$) time steps, each represented as a real number. We first rescale each number and represent it using $n$ digits (typically, $n=3$). Each time series is rescaled to the interval $[1.50,8.50]$ so that the number of digits never changes throughout the series. We then follow the scheme introduced in Gruver et al. (2023) to serialize the time series as strings and tokenize them.
@@ 3.2 Extraction of transition rules
Discrete state space.
When the Markov process is discrete and has a finite state space, each state can be represented by a single token. We employ tokens corresponding to the ten number strings: $0,\ldots,9$. We find that even the most sophisticated LLaMA model (LLaMA-70b) can only learn up to 9 discrete states. Therefore, we do not attempt to go beyond 9 distinct states by extending to non-number tokens (see Section A.3.3 for more details).
@@
Figure 2 illustrates our framework for learning discrete Markov chains with LLMs. First, we randomly sample an $n\times n$ transition matrix $(P_{ij})$. We then generate a Markov chain according to $P_{ij}$, tokenize the time series and pass to an LLM with no additional “prompt engineering". The length of the series is chosen such that the tokenized representation does not exceed the length of the LLM’s context window. We extract the LLM’s prediction for the next state by performing a softmax operation on the output logits corresponding to the $n$ allowed states and discarding all other logits.
Continuous state space.
Many stochastic processes, such as the Brownian motion Einstein (1905); Perrin (1909), are supported on continuous state space. For these processes, we represent the value of each state as a multi-digit number and separate each state using the comma symbol “,”. As observed in Jin et al. (2023a), an LLM prediction of multi-digit values can be naturally interpreted as a hierarchical softmax distribution Mnih and Hinton (2008); Challu et al. (2023).
@@
Hierarchy-PDF algorithm.
While a single pass through the LLM yields a discretized PDF represented by bins of various widths, we can refine the PDF by querying each coarse bin. For example, to furnish a maximal resolution PDF of a 3-digit number, we need to query all $10^{2}$ combinations of the first two digits of that number. Suppose a time series consists of $S$ values (steps), each represented as $n$ digits. Obtaining a maximal resolution PDF for each value of the entire sequence requires $10^{n-1}S$ forward passes of the LLM. This daunting process could be significantly simplified because most of the $10^{n-1}S$ inputs differ only in the last tokens, and thus one can recursively cache the key and value matrices associated with the shared tokens. The computation can be further reduced by refining only the high-probability bins near the mode, which dominate the loss functions, as shown in Figure 3. Algorithm 1 outlines the Hierarchy-PDF algorithm used to recursively refine the PDF associated with a multi-digit value in a time series (more details available in  Appendix A).
@@ 4 Experiments and Results
This section reports empirical in-context learning results on two example systems: a discrete Markov chain and a stochastic logistic map. We defer discussion of other systems, reported in Figure 1, to Sections A.3 and A.4. Each experiments is repeated ten times with trajectories initiated by different random seeds.
Model choice.
All numerical experiments reported in this section are performed using the open-source LLaMA-13b model. While we observe that larger language models, such as LLaMA-70b, may achieve lower in-context loss on some dynamical systems (Section A.3.3), they do not display qualitative differences that would affect our conclusions.
@@ 4.1 Markov chains with discrete states
Using the testing procedures elaborated on in Section 3.2, we generate 10 Markov chains, each from a distinct and randomly generated transition matrix of size $n=4$.
@@ 4.2 Noisy logistic map
The logistic map, first proposed as a discrete-time model for population growth, is one of the simplest dynamical systems that manifest chaotic behaviors Strogatz (2015). It is governed by the following iterative equation:
@@
where $r\in[1,4)$ is a parameter that controls the system’s behavior. [^3] In our experiments, we set $r=3.9$, which places the system firmly in the chaotic regime. At this value, the logistic map exhibits sensitive dependence on initial conditions and aperiodic behavior characteristic of chaos. The logistic map system becomes stochastic when one introduces small Gaussian perturbations of variance at each step, resulting in modified iterative equation:
@@
The first derivative of $f$ measures how sensitive the local dynamics are to external perturbations. This intuitively explains why the standard deviation of the conditional distribution should be proportional to $f^{\prime}$. [^4] Note that the approximation in Equation 1 assumes a small perturbation compared to the second derivative. That is, $\sigma^{2}\ll 1/f^{\prime\prime}(x)$. We visualize the remarkable accuracy of the LLM’s in-context predictions compared to the ground-truth transition rules in Figure 7 as well as Figure 19 in the appendix.
Figure 5: LLM’s prediction for the $980^{th}$ state of a noisy logistic map, conditioned on observing the first $979$ states in-context, compared with the ground-truth transition probability function.
@@
We again observe a power-law-like decay of the in-context loss function with respect to the length of the observed time series in Figure 7. We note that in order to achieve low in-context loss, the LLM must learn to predict not only the mean, but also the variance of next-state distributions. This is shown in Figure 7 and discussed further in Section A.9.
@@ 5 Discussion and conclusion
Main observations.
We showed that, with sufficient context, LLMs can accurately recover the probablistic transition rules underlying deterministic, chaotic, and stochastic time series.
In-context neural scaling law.
@@
Neural scaling laws Kaplan et al. (2020) are power laws that characterize how the loss of trained neural networks vary with respect to parameters of the model, such as model size, dataset size, and computational resources. To the best of our knowledge, neural scaling laws have so far only been observed in the training procedure, which updates the weights of neural networks using an explicit algorithm, such as stochastic gradient descent and Adam Kingma and Ba (2014). The loss curves observed in the different numerical experiments (see Figures 1 and 20) reveal an additional in-context scaling law for LLMs’ zero-shot learning of dynamical systems. Further analysis of these scaling laws are presented in Section A.6.
Figure 8: Loss curves of LLM in-context learning against baseline models: unigram and bi-gram models for discrete Markov chains, and linear and non-linear autoregressive models with 1-step memory (AR1) for noisy logistic maps. The coefficient $\alpha$ denotes the fitted scaling exponent.
Mechanistic interpretations of in-context learning.
Characterizing the in-context learning algorithms implemented by an LLM is an open question of broad interest  Shen et al. (2023); Olsson et al. (2022). Recent years have seen a new paradigm that explains certain aspects of in-context learning as transformers “hosting" classical algorithms such as gradient descent Von Oswald et al. (2023). Our study contributes to this field by presenting novel phenomenological evidence: we demonstrate that pretrained LLMs can in-context learn dynamical systems, exhibiting loss curves similar to those of established learning algorithms, such as the bi-gram Markov model and neural network-based autoregressive (AR1) models trained using gradient descent (Figure 8). We elaborate on the motivation and design of these baseline models in Appendix A.7.
Future directions.
A compelling future direction is to develop a mechanistic theory that explains how pretrained LLMs in-context learn and predict stochastic dynamical time series. Further investigations could focus on:
  • Investigating how LLMs encode spatio-temporal features in stochastic dynamical time series (Gurnee and Tegmark, 2023), and how they represent the learned transition rules.
  • Examining how LLMs perform subtasks necessary for learning stochastic dynamical systems, such as collecting Markovian transition statistics from discrete data (Akyürek et al., 2024) and performing density estimation.
  • Assessing to what extent training on natural language enhances a transformer’s mathematical ability (Zhang et al., 2023) to in-context learn dynamical systems, and vice-versa.
Limitations
Data Leakage.
Although the accurate predictions made by LLMs of the next time step value are highly unlikely to be due to memorization, it is essential to address this possibility thoroughly. Given the vast number of potential sequences even with a thousand numerical values encoded to three digits (resulting in $10^{3000}$ instance), this far exceeds the approximate $10^{12}$ tokens in the training corpus Touvron et al. (2023). Future work should ensure rigorous testing to further rule out data leakage.
Model selection.
At the time of this manuscript’s preparation, newer models, such as LLaMA-3, have been released. Due to constraints in time and computational resources, this work does not evaluate these latest models. Future research should include these and other emerging models to provide a more comprehensive understanding of the phenomena observed.
Acknowledgements
This work was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Numbers N00014-23-1-2729 and N00014-23-1-2716.
@@ A.1 Loss Functions
Bhattacharyya distance.
For stochastic time series, we use the Bhattacharyya distance to characterize the distance between learned and ground truth transition functions. The Bhattacharyya distance between $P$ and $\tilde{P}$, on a domain $\mathcal{X}$ is defined as Bhattacharyya (1943, 1946); Kailath (1967):
@@
and has been widely employed by feature selection and signal extraction methods Choi and Lee (2003); Kailath (1967). Since $\tilde{P}(X_{t+1}|X_{t})$ takes the form of a hierarchical PDF, one may approximate the integral in Equation 2 via a discrete quadrature rule as
@@
Squared deviations from the mean.
For deterministic systems, the true transition functions become delta-functions. As a result, the discretized Bhattacharyya distance from Equation 3 reduces to (see Equation 5 in Section A.2)
@@
which is proportional to the negative log-likelihood (NLL) assigned to the true data by the LLM, plus a constant $C$ [^5] This constant is determined by the base B of the system, and the number of digits n as $C=-\ln\Delta x=n\log B$.. NLL references only the finest bins in the hierarchical PDF and is thus unstable as an in-context loss. As an alternative, we use the squared deviations from the mean (SDM) Kobayashi and Salam (2000) as the in-context loss for deterministic systems:
@@ A.2 Additional loss functions
KL-divergence.
The KL-divergence between two PDFs, $P$ and $\tilde{P}$, is defined as
@@
Although commonly used as the training loss for a variety of machine learning systems, this loss function may suffer from numerical instabilities as the learned transition function $\tilde{P}$ are often close to zero, as shown in Figures 15 and 10, where the probability density is concentrated in small regions of the support.
Discretized Bhattacharyya distance for deterministic systems.
For deterministic systems, the ground truth transition function is a delta function. Therefore, the Bhattacharyya distance between it and the Hierarchy-PDF prediction only references the finest bin associated with the true value $x_{\mathrm{true}}$.
@@
Higher moments and kurtosis.
While the Bhattacharyya distance and SDM measure the agreement between the extracted transition rules $\tilde{P}$ and the ground truth distribution $P$, they do not explicitly characterize the type of the distribution (e.g., Gaussian or uniform). We employ the kurtosis as an additional measure to assess whether the LLM recovers the correct shape of $P$. The kurtosis of a distribution $P$ is defined as Joanes and Gill (1998)
@@ A.3.1 Brownian motion
Brownian motion is an example of a continuous-time stochastic process Einstein (1905), and is described by a stochastic differential equation (SDE):
@@
where $X_{t}$ represents the state of the system at time $t$, $\mu$ is the drift coefficient, $\sigma$ is the volatility coefficient, and $dW_{t}$ is the increments of a Wiener process Revuz and Yor (2013), modeling the randomness of motion.
@@
To simulate trajectories of Brownian motion, we use the Euler–Maruyama method Platen (1999), which discretizes Equation 8 as $X_{t+\Delta t}=X_{t}+\mu\Delta t+\sigma\sqrt{\Delta t}Z$, where $\Delta t$ is the time resolution, and $Z\sim\mathcal{N}(0,1)$ is a random variable that follows a standard Gaussian distribution. The Euler–Maruyama method may also be written as a conditional distribution: